# C-LoRA: Contextual Low-Rank Adaptation for Uncertainty Estimation in Large Language Models

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) offers a cost-effective solution for fine-tuning large language models (LLMs), but it often produces overconfident predictions in data-scarce few-shot settings. To address this issue, several classical statistical learning approaches have been repurposed for scalable uncertainty-aware LoRA fine-tuning. However, these approaches neglect how input characteristics affect the predictive uncertainty estimates. To address this limitation, we propose Contextual Low-Rank Adaptation (C-LoRA) as a novel uncertainty-aware and parameter efficient fine-tuning approach, by developing new lightweight LoRA modules contextualized to each input data sample to dynamically adapt uncertainty estimates. Incorporating data-driven contexts into the parameter posteriors, C-LoRA mitigates overfitting, achieves well-calibrated uncertainties, and yields robust predictions. Extensive experiments demonstrate that C-LoRA consistently outperforms the state-of-the-art uncertainty-aware LoRA methods in both uncertainty quantification and model generalization. Ablation studies further confirm the critical role of our contextual modules in capturing sample-specific uncertainties. C-LoRA sets a new standard for robust, uncertainty-aware LLM fine-tuning in few-shot regimes.
@@ 1 Introduction
Large Language Models (LLMs) (shorinwa2024surveyuncertaintyquantificationlarge, ; minaee2024largelanguagemodelssurvey, ; radford2019language, ; brown2020languagemodelsfewshotlearners, ; openai2024gpt4technicalreport, ; touvron2023llama2openfoundation, ; grattafiori2024llama3herdmodels, ) have shown their promising potential in diverse areas (Cui2024, ; Luo_2022, ; Singhal2023-zh, ; yang2022gatortronlargeclinicallanguage, ; jantre2025uqllm, ; bran2023chemcrowaugmentinglargelanguagemodels, ; taylor2022galacticalargelanguagemodel, ; yang2023fingptopensourcefinanciallarge, ; wu2023bloomberggptlargelanguagemodel, ; thulke2024climategptaisynthesizinginterdisciplinary, ; chen2021evaluatinglargelanguagemodels, ). Due to their general-purpose language understanding and generation capabilities with human-level performance (shorinwa2024surveyuncertaintyquantificationlarge, ; minaee2024largelanguagemodelssurvey, ), fine-tuning LLMs to various downstream tasks has drawn significant attention (yang2024bayesianlowrankadaptationlarge, ; wang2024blobbayesianlowrankadaptation, ; hu2021loralowrankadaptationlarge, ; lialin2024scalingscaleupguide, ; wang2022listlitepromptedselftraining, ). However, when fine-tuned for downstream tasks with limited data, LLMs may hallucinate to produce overconfident results, which becomes a serious concern (li2024thinktwicetrustingselfdetection, ; xiong2024llmsexpressuncertaintyempirical, ; leng2024tamingoverconfidencellmsreward, ; he2023preservingpretrainedfeatureshelps, ). To address this, reliable estimation of uncertainty has become essential wang2024blobbayesianlowrankadaptation ; yang2024bayesianlowrankadaptationlarge ; meo2024bayesianloralorabasedparameter .
@@
To enable predictive uncertainty quantification (UQ), probabilistic inference with Bayesian neural network (BNN) has been studied in deep learning pmlr-v48-gal16 ; zhou2019bayesianinferenceuncertaintyquantification ; shaker2021ensemblebaseduncertaintyquantificationbayesian ; wang2024uncertaintyquantificationnearbayesoptimal ; Goan_2020 ; daxberger2022laplacereduxeffortless ; zhang2021bayesianattentionbeliefnetworks ; zhang2020cyclicalstochasticgradientmcmc ; kristiadi2020bayesianjustbitfixes ; kendall2017uncertaintiesneedbayesiandeep ; fan2021contextualdropoutefficientsampledependent ; jantre2022sequential ; jantre2024active , where neural network weights are treated as random variables with variational inference (VI) approximating the true posterior to provide reliable UQ (Blei03042017, ; hoffman2013stochasticvariationalinference, ). Adopting these approaches for LLMs is limited by their prohibitive computational and memory costs compared to conventional methods yang2024bayesianlowrankadaptationlarge ; wang2024blobbayesianlowrankadaptation ; lialin2024scalingscaleupguide , making full Bayesian fine-tuning of all the parameters of an LLM challenging.
@@
Parameter Efficient Fine-Tuning (PEFT) methods, such as Low-rank Adaptation (LoRA), substantially reduce the number of learnable parameters and thereby mitigate the significant computational expenses and excessive memory utilization hu2021loralowrankadaptationlarge ; Ding2023 ; liu2022fewshotparameterefficientfinetuningbetter ; shi2024deptdecomposedprompttuning ; han2024parameterefficientfinetuninglargemodels ; edalati2022kronaparameterefficienttuning . This efficiency facilitates employing Bayesian methods for UQ in LLMs balabanov2024uncertaintyquantificationfinetunedllms ; onal2024gaussianstochasticweightaveraging ; yang2024bayesianlowrankadaptationlarge ; wang2024blobbayesianlowrankadaptation ; wang2023loraensembleslargelanguage . Most of the recent studies considered a more straightforward solution, like in balabanov2024uncertaintyquantificationfinetunedllms ; onal2024gaussianstochasticweightaveraging ; wang2023loraensembleslargelanguage where authors considered ensemble methods. yang2024bayesianlowrankadaptationlarge proposed a post-hoc Bayesian inference method for the LoRA adapters using Laplace approximation daxberger2022laplacereduxeffortless . wang2024blobbayesianlowrankadaptation repurposed a mean-field VI based Bayesian framework for LoRA-based fine-tuning to jointly estimate the LoRA parameters’ variational means and variances wang2024blobbayesianlowrankadaptation . However, the uncertainty stemming from data, aleatoric uncertainty, has not been considered in the existing methods, leading to poor performance especially when fine-tuning with limited data. This motivates us to focus on incorporating aleatoric uncertainty for a novel parameter efficient fine-tuning approach in this work.
@@
    Through ablation experiments, we demonstrate the significance of our new contextual module on achieving better UQ while offering consistent prediction accuracy.
@@ 2 Preliminaries
In this paper, vectors and matrices are denoted by bold lowercase and uppercase letters, respectively. Most of our mathematical notations follow the ones adopted in hu2021loralowrankadaptationlarge , yang2024bayesianlowrankadaptationlarge , wang2024blobbayesianlowrankadaptation , and fan2021contextualdropoutefficientsampledependent .
@@ 2.1 Low-Rank Adaptation (LoRA)
LoRA, a parameter-efficient fine-tuning approach, adapts a pre-trained language model to downstream tasks hu2021loralowrankadaptationlarge . It rests on the assumption that the required weight updates have a low intrinsic dimensionality, so LoRA freezes the original weights and instead learns low-rank update matrices. To this end, the modified forward pass becomes:
@@ 2.2 Bayesian Uncertainty Estimation
[…]
@@ 3 Contextual Low-Rank Adaptation (C-LoRA)
We introduce the formulation of Contextual Low-Rank Adaptation (C-LoRA), which enables scalable, efficient, data-dependent uncertainty quantification at the sample level. First, we introduce a lightweight LoRA factorization to reduce the computational burden of variational inference in Bayesian LoRA. Based on this factorization, we treat the LoRA weights’ stochasticity to be data-dependent, and learn the weight distribution parameters via a variational Bayesian objective.
@@ 3.1 Lightweight LoRA Factorization
The standard LoRA update in (1) introduces two low-rank matrices whose size scales with the frozen weight dimension $d$, so that any Bayesian treatment—whether both adapters are stochastic or only $\mathbf{A}$ is (as in BLoB)—incurs a computational cost that grows linearly in $d$. To break this dependency, we insert a low-dimensional matrix $\mathbf{E}\in\mathbb{R}^{r\times r}$ between $\mathbf{B}$ and $\mathbf{A}$. This modification reduces the stochastic parameter complexity to a constant in $d$, yielding a modified LoRA factorization as follows
@@
Compared to the BLoB mean-field VI framework, this new factorization facilitates a more scalable, lightweight Bayesian LoRA by inferring a distribution over the elements of $\mathbf{E}$ elements while learning $\mathbf{B}$ and $\mathbf{A}$ deterministically, making data-dependent Bayesianization of LoRA fine-tuning scalable.
3.2 Contextualized Parameter Uncertainty
In conventional Bayesian neural networks, the focus is on epistemic uncertainties introduced by treating model parameters as random variables drawn from a fixed distribution—one that does not depend on the input data. This leads to model parameters that can vary from data to data, while their distribution remains invariant for all the samples within given training data. This severely limits the uncertainty estimates in few-shot LoRA fine-tuning settings. To address this, we propose a data-dependent, or contextual, Bayesian fine-tuning paradigm, wherein the distribution of the parameters of low-rank adapters depends on the inputs $\mathbf{x}_{i}$ for each data sample $(\mathbf{x}_{i},y_{i})$. Although one could perform Bayesian inference over $\mathbf{B}$ and $\mathbf{A}$ LoRA adapters from (1), this scales linearly with the frozen weight dimension $d$. Instead, by leveraging the lightweight LoRA factorization from (4), we learn the low-dimensional $\mathbf{E}$ matrix contextually, yielding the weight updates as
@@ 3.2 Contextualized Parameter Uncertainty
[…]
$\mathcal{L}=\sum_{i=1}^{N}\left[\mathbb{E}_{\,q_{\phi}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})}\log p_{\boldsymbol{\theta}}(y_{i}|\mathbf{x}_{i},\mathbf{E}_{\mathbf{x}_{i}})-\mathrm{KL}(q_{\phi}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i}))\,\|\,p(\mathbf{E}_{\mathbf{x}_{i}}))\right].$ (6)
Here, $\boldsymbol{\theta}$ represents $\mathbf{B}$ and $\mathbf{A}$ collectively. Unlike standard variational inference as in (3), our formulation makes the distribution of $\mathbf{E}_{\mathbf{x}_{i}}$ input-dependent and replaces the single KL term with a sum of $N$ per-sample $\mathrm{KL}$ divergences, whose combined impact scales with $N$. More importantly, we focus exclusively on aleatoric uncertainty—modeling the variability in $y_{i}$ given $\mathbf{x}_{i}$—with the epistemic uncertainty modeling available via imposing prior on $\varphi$ or Bayesianizing $\mathbf{A}$ or $\mathbf{B}$ when needed, thereby isolating and highlighting the advantages of data-dependent aleatoric UQ.
Sequential Cross-Layer Dependency. Consider a pre-trained LLM with $L$ layers, each augmented by its own LoRA adapters $\mathbf{B}^{l},\mathbf{E}_{\mathbf{x}}^{l}\,,$ and $\mathbf{A}^{l}$. We model $\mathbf{E}_{\mathbf{x}_{i}}^{l}$ in an autoregressive manner, so that its distribution depends on $\{\mathbf{E}_{\mathbf{x}_{i}}^{j}\}_{j<l}$, leading to $q_{\phi}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})=\prod_{l=1}^{L}q_{\phi}(\mathbf{E}_{\mathbf{x}_{i}}^{l}|\mathbf{x}_{i}^{l-1})$, where $\mathbf{x}_{i}^{l-1}$ denotes the output of layer $l-1$ and is a function of $\{\mathbf{E}_{\mathbf{x}_{i}}^{1},\cdots,\mathbf{E}_{\mathbf{x}_{i}}^{l-1},\mathbf{x}_{i}\}$.
3.3 Auxiliary Contextual Module Parameterization
@@
In layer $l$, given $\mathbf{z}^{l}=\mathbf{A}^{l}\mathbf{x}^{l-1}\in\mathbb{R}^{r}$ as input, the contextual module denoted by $h^{l}_{\varphi}$ produces the parameters $(\boldsymbol{\mu}^{l}_{\mathbf{E}},\boldsymbol{\Omega}^{l}_{\mathbf{E}})$ of the $\mathbf{E}^{l}_{\mathbf{x}}$ distribution as output. We then draw $\mathbf{E}^{l}_{\mathbf{x}}$ conditioned on $\boldsymbol{\mu}^{l}_{\mathbf{E}}$ and $\boldsymbol{\Omega}^{l}_{\mathbf{E}}$. Then, multiplying $\mathbf{B}^{l}$, $\mathbf{E}^{l}_{\mathbf{x}}$, and $\mathbf{z}^{l}$ and adding it to $\mathbf{W}^{l}_{0}\,\mathbf{x}^{l-1}$ yields the output of the layer $l$, $\mathbf{x}^{l}$. We parameterize the per-layer contextual module with a neural network with two fully-connected layers, parametrized by $\varphi$. In particular, the neural network has $r$ inputs, $C$ hidden units, and $2\times r^{2}$ outputs with the nonlinear ReLU activation function connecting the two fully-connected layers.
3.4 Variational Bayesian Objective
Our optimization objective for C-LoRA fine-tuning is as given in (6). We adopt a simple and fixed Gaussian prior $p(\mathbf{E_{x}})$ for learning $\mathbf{E_{x}}$ in each layer. The complete learning objective can be expressed as a summation over all samples: $\mathcal{L}=\sum_{(\mathbf{x},y)\in\mathcal{D}}\mathcal{L}(\mathbf{x},y)$. Then, we learn the deterministic parameters $\mathbf{B}$ and $\mathbf{A}$ together represented by $\boldsymbol{\theta}$ using the expected negative log-likelihood (NLL) term while excluding the $\mathrm{KL}$ term. This ensures that learning $\boldsymbol{\theta}$ remains supervised, yielding less noisy gradients per sample computed as
$\nabla_{\boldsymbol{\theta}}\mathcal{L}(\mathbf{x},y)=\mathbb{E}_{\,q_{\phi}(\mathbf{E}_{\mathbf{x}}|\mathbf{x})}\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(y|\mathbf{x},\mathbf{E}_{\mathbf{x}}).$ (7)
We approximate this expectation with a single Monte Carlo draw $\mathbf{E_{x}}\sim q_{\phi}(\mathbf{E_{x}}|\mathbf{x})$ for each $\mathbf{x}$ sample.
Next, to update the auxiliary network parameters $\varphi$, which appear in both the NLL and $\mathrm{KL}$ terms of (6), we apply the reparameterization trick kingma2015repara_trick to handle random sampling of $\mathbf{E_{x}}$. Concretely, in each layer, $\mathbf{E}^{l}_{\mathbf{x}}\sim\mathcal{N}(\boldsymbol{\mu}^{l}_{\mathbf{E}},{\boldsymbol{\Omega}^{l}}^{2}_{\mathbf{E}})$ is reparameterized as $\mathbf{E}^{l}_{\mathbf{x}}=\boldsymbol{\mu}^{l}_{\mathbf{E}}+\boldsymbol{\Omega}^{l}_{\mathbf{E}}\odot\mathcal{E}^{l}$, where $\mathcal{E}^{l}\in\mathbb{R}^{r\times r}$ is sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Similarly, sampling the sequence of $\mathbf{E}^{l}_{\mathbf{x}}$ for all the layers using the distribution $q_{\phi}(\mathbf{E}_{\mathbf{x}}|\mathbf{x})$ is rewritten as $g_{\phi}(\mathcal{E},\mathbf{x})$, where $g_{\phi}$ is a deterministic and differentiable mapping. With this, the gradient of the fine-tuning objective $\mathcal{L}$ with respect to $\varphi$ can now be computed as

### v3

@@ Abstract
Low-Rank Adaptation (LoRA) offers a cost-effective solution for fine-tuning large language models (LLMs), but it often produces overconfident predictions in data-scarce few-shot settings. To address this issue, several classical statistical learning approaches have been repurposed for scalable uncertainty-aware LoRA fine-tuning. However, these approaches neglect how input characteristics affect the predictive uncertainty estimates. To address this limitation, we propose Contextual Low-Rank Adaptation (C-LoRA) as a novel uncertainty-aware and parameter efficient fine-tuning approach, by developing new lightweight LoRA modules contextualized to each input data sample to dynamically adapt uncertainty estimates. Incorporating data-driven contexts into the parameter posteriors, C-LoRA mitigates overfitting, achieves well-calibrated uncertainties, and yields robust predictions. Extensive experiments on LLaMA2-7B models demonstrate that C-LoRA consistently outperforms the state-of-the-art uncertainty-aware LoRA methods in both uncertainty quantification and model generalization. Ablation studies further confirm the critical role of our contextual modules in capturing sample-specific uncertainties. C-LoRA sets a new standard for robust, uncertainty-aware LLM fine-tuning in few-shot regimes. Although our experiments are limited to 7B models, our method is architecture-agnostic and, in principle, applies beyond this scale; studying its scaling to larger models remains an open problem. Our code is available at https://github.com/ahra99/c_lora.
@@ 1 Introduction
To enable predictive uncertainty quantification (UQ), probabilistic inference with Bayesian neural network (BNN) has been studied in deep learning pmlr-v48-gal16 ; doi:10.1137/19M1248352 ; Shaker-2021 ; wang2024on ; Goan_2020 ; NEURIPS2021_a7c95857 ; pmlr-v139-zhang21f ; Zhang2020Cyclical ; pmlr-v119-kristiadi20a ; NIPS2017_2650d608 ; fan2021contextual ; jantre2022sequential ; jantre2024active , where neural network weights are treated as random variables with variational inference (VI) approximating the true posterior to provide reliable UQ (Blei03042017, ; JMLR:v14:hoffman13a, ). Adopting these approaches for LLMs is limited by their prohibitive computational and memory costs compared to conventional methods yang2024bayesianlowrankadaptationlarge ; wang2024blobbayesianlowrankadaptation ; lialin2024scalingscaleupguide , making full Bayesian fine-tuning of all the parameters of an LLM challenging.
@@
Parameter Efficient Fine-Tuning (PEFT) methods, such as Low-rank Adaptation (LoRA), substantially reduce the number of learnable parameters and thereby mitigate the significant computational expenses and excessive memory utilization hu2022lora ; Ding2023 ; liu2022fewshot ; shi2024dept ; han2024parameterefficient ; Edalati2025 . This efficiency facilitates employing Bayesian methods for UQ in LLMs balabanov2025uncertainty ; onal2024gaussian ; yang2024bayesianlowrankadaptationlarge ; wang2024blobbayesianlowrankadaptation ; wang2023loraensembleslargelanguage . Most of the recent studies considered a more straightforward solution, like in balabanov2025uncertainty ; onal2024gaussian ; wang2023loraensembleslargelanguage where authors considered ensemble methods. yang2024bayesianlowrankadaptationlarge proposed a post-hoc Bayesian inference method for the LoRA adapters using Laplace approximation NEURIPS2021_a7c95857 . wang2024blobbayesianlowrankadaptation repurposed a mean-field VI based Bayesian framework for LoRA-based fine-tuning to jointly estimate the LoRA parameters’ variational means and variances wang2024blobbayesianlowrankadaptation . However, the uncertainty stemming from data, aleatoric uncertainty, has not been considered in the existing methods, leading to poor performance especially when fine-tuning with limited data. This motivates us to focus on incorporating aleatoric uncertainty for a novel parameter efficient fine-tuning approach in this work.
@@
    Through ablation experiments, we demonstrate the significance of our new contextual module on achieving better UQ while offering competitive accuracy with only minor drops on some tasks.
@@ 2 Preliminaries
In this paper, vectors and matrices are denoted by bold lowercase and uppercase letters, respectively. Most of our mathematical notations follow the ones adopted in hu2022lora , yang2024bayesianlowrankadaptationlarge , wang2024blobbayesianlowrankadaptation , and fan2021contextual .
@@ 2.1 Low-Rank Adaptation (LoRA)
LoRA, a parameter-efficient fine-tuning approach, adapts a pre-trained language model to downstream tasks hu2022lora . It rests on the assumption that the required weight updates have a low intrinsic dimensionality, so LoRA freezes the original weights and instead learns low-rank update matrices. To this end, the modified forward pass becomes:
@@ 2.2 Bayesian Uncertainty Estimation
$p(y^{*}|\mathbf{x}^{*},\mathcal{D})=\int p(y^{*}|\mathbf{x}^{*},\boldsymbol{\theta})\,p(\boldsymbol{\theta}|\mathcal{D})\,d\boldsymbol{\theta}\;\approx\frac{1}{M}\sum_{m=1}^{M}p(y^{*}|\mathbf{x}^{*},\boldsymbol{\theta}_{m}),\quad\boldsymbol{\theta}_{m}\sim p(\boldsymbol{\theta}|\mathcal{D}).$ (2)
@@ 3 Contextual Low-Rank Adaptation (C-LoRA)
We introduce the formulation of Contextual Low-Rank Adaptation (C-LoRA), which enables scalable, efficient, data-dependent uncertainty quantification at the sample level. First, we introduce a lightweight LoRA factorization to reduce the computational burden of variational inference in Bayesian LoRA. Based on this factorization, we treat the LoRA weights’ stochasticity to be data-dependent, and learn the weight distribution parameters via a variational Bayesian objective. A step-by-step description of C-LoRA is presented in Algorithm 1 (Appendix F).
@@ 3.1 Lightweight LoRA Factorization
The standard LoRA update in (1) introduces two low‐rank matrices whose size scales with the frozen weight dimension $d$, so that any Bayesian treatment—whether both adapters are stochastic or only $\mathbf{A}$ is (as in BLoB)—incurs a computational cost that grows linearly in $d$. To break this dependency, we insert a low-dimensional matrix $\mathbf{E}\in\mathbb{R}^{r\times r}$ between $\mathbf{B}$ and $\mathbf{A}$. This modification reduces the stochastic parameter complexity to a constant in $d$, yielding a modified LoRA factorization as follows
@@
Compared to the BLoB mean‐field VI framework, this new factorization facilitates a more scalable, lightweight Bayesian LoRA by inferring a distribution over the elements of $\mathbf{E}$ while learning $\mathbf{B}$ and $\mathbf{A}$ deterministically, making data-dependent Bayesianization of LoRA fine-tuning scalable.
3.2 Stochastic LoRA Parameterization with Data Dependence
In conventional Bayesian neural networks, the focus is on epistemic uncertainties introduced by treating model parameters as random variables drawn from a fixed distribution—one that does not depend on the input data. Consequently, while the sampled parameters may vary across data instances, their underlying distribution remains invariant across all samples in the training set. This assumption limits the expressiveness of uncertainty estimates, particularly in few-shot LoRA fine-tuning scenarios. To address this, we propose a data-dependent, or contextual, Bayesian fine-tuning paradigm, wherein the distribution of the parameters of low-rank adapters depends on the inputs $\mathbf{x}_{i}$ for each data sample $(\mathbf{x}_{i},y_{i})$. Although one could perform Bayesian inference over $\mathbf{B}$ and $\mathbf{A}$ LoRA adapters from (1), this scales linearly with the frozen weight dimension $d$. Instead, by leveraging the lightweight LoRA factorization from (4), we learn the low-dimensional $\mathbf{E}$ matrix contextually, yielding the weight updates as
@@ 3.2 Stochastic LoRA Parameterization with Data Dependence
where $\mathbf{E_{x}}$ denotes the data-dependent version of $\mathbf{E}$. To be specific, we take an input-dependent Gaussian variational posterior over $\mathbf{E}_{\mathbf{x}_{i}}$, $q_{\boldsymbol{\phi}}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})=\mathcal{N}(\boldsymbol{\mu}_{\mathbf{E}}(\mathbf{x}_{i}),\boldsymbol{\Omega}^{2}_{\mathbf{E}}(\mathbf{x}_{i}))$. We learn these distribution parameters via lightweight per-layer auxiliary contextual modules comprising of small neural networks whose parameters across modules are collectively denoted as $\boldsymbol{\varphi}$. Such an input-dependent variational posterior resembles amortized inference in Bayesian modeling where $q_{\boldsymbol{\phi}}$ serves as an inference network that approximates $p(\mathbf{E}_{\mathbf{x}_{i}}|y_{i},\mathbf{x}_{i})\propto p(y_{i}|\mathbf{x}_{i},\mathbf{E}_{\mathbf{x}_{i}})p(\mathbf{E}_{\mathbf{x}_{i}})$.
@@
Particularly, consider a pre-trained LLM with $L$ layers, each augmented by its own LoRA adapters $\mathbf{B}^{l},\mathbf{E}_{\mathbf{x}}^{l}\,$, and $\mathbf{A}^{l}$. Let $\mathbf{x}_{i}^{l-1}$ denote the output of layer l1l-1. We then model $\mathbf{E}_{\mathbf{x}_{i}}^{l}$ autoregressively by conditioning on $\{\mathbf{E}_{\mathbf{x}_{i}}^{j}\}_{j<l}$, leading to $q_{\phi}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})=\prod_{l=1}^{L}q_{\phi}(\mathbf{E}_{\mathbf{x}_{i}}^{l}|\mathbf{x}_{i}^{l-1})$. Especially, in layer $l$, given $\mathbf{z}^{l}=\mathbf{A}^{l}\mathbf{x}^{l-1}\in\mathbb{R}^{r}$ as input, the corresponding contextual module denoted by $h^{l}_{\varphi}$ produces the parameters $(\boldsymbol{\mu}^{l}_{\mathbf{E}},\boldsymbol{\Omega}^{l}_{\mathbf{E}})$ of the $\mathbf{E}^{l}_{\mathbf{x}}$ distribution as output. We then draw $\mathbf{E}^{l}_{\mathbf{x}}$ conditioned on $\boldsymbol{\mu}^{l}_{\mathbf{E}}$ and $\boldsymbol{\Omega}^{l}_{\mathbf{E}}$. Finally, multiplying $\mathbf{B}^{l}$, $\mathbf{E}^{l}_{\mathbf{x}}$, and $\mathbf{z}^{l}$ and adding it to $\mathbf{W}^{l}_{0}\,\mathbf{x}^{l-1}$ yields the output of the layer $l$, $\mathbf{x}^{l}$.
This formulation enables us to focus exclusively on heteroscedastic uncertainty—modeling the variability in $y_{i}$ given $\mathbf{x}_{i}$—with the epistemic uncertainty modeling available via imposing prior on $\boldsymbol{\varphi}$ or Bayesianizing $\mathbf{A}$ or $\mathbf{B}$ when needed, thereby isolating and highlighting the advantages of data‐dependent (heterscedetic) UQ.
Contextual module parameterization. We parameterize each layer’s contextual module with a two fully-connected layer neural network whose parameters at layer $l$ are denoted as $\varphi^{l}$. In particular, each of these neural networks have $r$ inputs, $C$ hidden units, and $2\times r^{2}$ outputs with the nonlinear ReLU activation function connecting the two fully-connected layers.
3.3 Amortized Inference for Contextual LoRA
In our formulation, the variational distribution is only conditioned on $\mathbf{x}_{i}$, while $y_{i}$ contributes through the training objective. We then learn the model parameters, $\boldsymbol{\phi}$, by maximizing the ELBO of $\sum_{i}\log p(y_{i}|\mathbf{x}_{i})=\sum_{i}\log\int p(y_{i}|\mathbf{x}_{i},\mathbf{E}_{\mathbf{x}_{i}})p(\mathbf{E}_{\mathbf{x}_{i}})d\mathbf{E}_{\mathbf{x}_{i}}$ given $q_{\boldsymbol{\phi}}(\mathbf{E}_{\mathbf{x}_{i}}|\mathbf{x}_{i})$: